13. (5 points) parametrization Find an equation of the tangent line to the curve given by the =t- - y = 1+12 at t=1.
Answer #1

(x, y) = (0,21 x= t-1 ti ti ya lt to at 11 t-1 x=1-1=0, y = 1 + 1 = 2 dy at=2 odt da 1+1 t=1 tr dt t=1 - 2 ma dy dx dy dt dt